OpenStudy (anonymous):

Determine where f(x) intersects the x-axis. f(x)=5x^2 +40x+80

OpenStudy (anonymous):

That, like, intersects the x axis at f(x)=0, so you should, like, 0=x^2+8x+16, or 0=(x+4)^2, meaning x=-4.

OpenStudy (anonymous):

Whoa, visuals!